4^(×)+4^(×+1)+3=0

Simplify


1
Expert's answer
2021-07-23T15:42:01-0400

"4^{\left(x\right)}+4^{\left(x+1\right)}+3=0\\
4^{\left(x+1\right)+x}+3=0\\
4^{2x+1}=-3\\
4^{2x}=-\frac{3}{4}\\
a^{f\left(x\right)}\mathrm{\:cannot\:be\:zero\:or\:negative\:for\:}x\in \mathbb{R}\\
\mathrm{No\:Solution\:for}\:x\in \mathbb{R}"